40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 24 21:02:46 2009, Alberto, Update, PSL, Aligning the beam to the Faraday 
    Reply  Sat Jul 25 16:04:32 2009, rob, Update, PSL, Aligning the beam to the Faraday 
       Reply  Sat Jul 25 19:04:01 2009, Koji, Update, PSL, Aligning the beam to the Faraday MC2_Cam.jpg
          Reply  Sun Jul 26 13:19:54 2009, rana, Update, PSL, Aligning the mode cleaner Untitled.png
             Reply  Sun Jul 26 16:05:17 2009, Alberto, Update, PSL, Aligning the mode cleaner 
             Reply  Mon Jul 27 09:34:07 2009, steve, Summary, IOO, Aligning the mode cleaner 20dtrend.jpg
Message ID: 1795     Entry time: Mon Jul 27 09:34:07 2009     In reply to: 1793
Author: steve 
Type: Summary 
Category: IOO 
Subject: Aligning the mode cleaner 

Quote:

I set the MC back to its good alignment (June 21st) using this procedure. The trend of the OSEM values over the last 40 days and 40 nights is attached.

Then I aligned the periscope to that beam. This took some serious periscope knob action. Without WFS, the transmission went to 2.7 V and the reflection down to 0.6V.

Then I re-aligned the MC_REFL path as usual. The beam was far enough off that I had to also re-align onto the MC LSC PD as well as the MC REFL camera (~2 beam radii).

Beams are now close to their historical positions on Faraday and MC2. I then restored the PZT sliders to their April snapshot and the X-arm locked.

Steve - please recenter the iris which is on the periscope. It has been way off for a long time.

So it looks OK now. The main point here is that we can trust the MC OSEMs.

Afterwards I rebooted c1susvme1 and c1susvme2 because they were skewed.

 

 I'm impressed by Rana's simple way to align the MC. IFO arms are locked or flashing. 20 days trend attached.

 

Attachment 1: 20dtrend.jpg  109 kB  | Show | Show all
ELOG V3.1.3-